# From Bugs to Benchmarks: A Comprehensive Survey of Software Defect Datasets

### v1

@@ 3.1.2. Defects in Compilers and Runtime Engines
Compilers and runtime engines are critical components of software systems, responsible for translating and executing code. Defects in compilers and runtime engines can significantly impact software performance, security, and execution correctness. We here mean by “runtime engine” both stand-alone implementations of a programming language, e.g., the Java Virtual Machine, and engines integrated into a larger software, e.g., a JavaScript engine embedded into a web browser. Datasets of defects in compilers and runtime engines often come as by-products of empirical research. For example, Yu and Wang (2024) conduct an empirical study on bugs in implementations of the Rust language (rus, 2025), presenting a dataset of 194 bugs from Rust compilers and Rust infrastructure. Xu et al. (2023) investigate security bugs introduced by compilers for C/C++, highlighting the potential risks when silent bugs go undetected. In addition, as also mentioned above, Du et al. (2021) and Shen et al. (2021) study defects in deep-learning compilers, identifying common root causes and symptoms. For runtime engines, Bugs in Pods (Yu et al., 2024) provides 429 bugs in the container runtime systems containerd (con, 2025), cri-o (cri, 2025), gvisor (gvi, 2025), and runc (run, 2025). Additionally, Paul et al. (2021) provide a detailed dataset of security-related changes in the Chromium OS project, and Zaman et al. (2012) present a dataset of performance-related bugs in web browsers, focusing on how such bugs degrade system performance. Besides browsers, datasets constructed from studies of other runtime engines include defects in WebAssembly engines (Wang et al., 2023c; Zhang et al., 2024a), and bugs that impact the correctness and performance of JavaScript engines (Wang et al., 2023a). Altogether, these datasets are critical for understanding and addressing defects in compilers and runtime environments.
@@ 3.1.3. Defects in Mobile Applications
Mobile applications are prone to various types of defects, including functional bugs, security vulnerabilities, and resource issues. Several datasets, mostly based on Android applications, have been developed to represent these defects. For example, LVDAndro (Senanayake et al., 2023) and Ghera (Mitra and Ranganath, 2017) focus on security vulnerabilities, providing labelled vulnerability datasets and benchmarks for AI-based detection models and vulnerability analysis. Functional bugs are collected by Andror2 (Wendland et al., 2021) and by Xiong et al. (2023), both of which analyze the behavior of defective mobile applications by manually reproducing the bugs. More specialized defects, such as data loss issues and resource leaks, are captured by DataLossRepository (Riganelli et al., 2019) and DroidLeaks (Liu et al., 2019), respectively. Finally, DroixBench (Tan et al., 2018) provides defective APK files particularly for evaluating automated repair techniques on Android applications. These datasets play a crucial role in enhancing the reliability and security of Android applications by supporting bug detection, repair, and vulnerability research.
@@ 3.1.4. Defects Related to Blockchain
Blockchain, as a decentralized and distributed ledger technology, has been widely used in various applications, such as cryptocurrencies (Zheng et al., 2017). In general, blockchain applications are implemented as smart contracts, which are self-executing contracts with the terms of the agreement between multiple parties directly written into code. Defects in blockchain applications can have severe implications, such as financial loss. Several datasets focus on identifying and analyzing defects in smart contracts. From a general perspective, Jiuzhou (Zhang et al., 2020) is a dataset of Ethereum smart contract bugs that are classified into 49 categories. Specifically for security vulnerabilities, SBP-Dataset (Cai et al., 2024) contains smart contract defects generated by removing guards and turning safe functions into unsafe ones. In addition, there are two datasets constructed from empirical studies. One defines 20 categories of smart contract bugs (Chen et al., 2022), and the other identifies common patterns of bug fixes (Wang et al., 2023b). Unlike the above datasets, Kim et al. (2022) focus on energy bugs in Ethereum client software instead of smart contracts. The dataset contains 507 bugs that may cause energy inefficiency mined from the buggy commits of client software operated in the Ethereum network.
@@ 3.1.5. Defects in Autonomous Systems
Autonomous systems, such as autonomous vehicles, aircraft, or robots, rely on robust software to operate safely and efficiently. Defects in autonomous systems can lead to severe consequences, such as accidents, injuries, and fatalities. Wang et al. (2021a) and Garcia et al. (2020) both present empirical studies of software defects in autonomous systems. The former study has collected a dataset of 569 bugs from PX4 (px4, 2025) and ArduPilot (ard, 2025), which are well-known open-source autopilot software systems for unmanned aerial vehicles. The latter study presents a dataset of 499 autonomous vehicle bugs from Apollo (apo, 2025) and Autoware (aut, 2025). In addition, ROBUST (Timperley et al., 2024) offers a dataset of 221 bugs from seven popular robot operating systems. Together, these datasets highlight the unique challenges caused by defects in autonomous systems.
@@ 3.1.6. Defects Related to Quantum Computing
Defects in quantum computing systems, including both quantum applications and quantum platforms, present unique challenges due to the complexity and novelty of the technology. Bugs4Q (Zhao et al., 2023a) provides a benchmark of bugs in quantum programs, particularly focusing on the Qiskit framework (qis, 2025), enabling controlled testing and debugging studies. Additionally, there are defect datasets constructed within empirical studies, such as one that gathers and categorizes bug patterns across various quantum platforms (Paltenghi and Pradel, 2022), and another that focuses on defects within the intersection of quantum computing and machine learning (Zhao et al., 2023b). These datasets are useful for advancing the reliability of quantum software systems.
@@ 3.1.7. Defects Related to Scientific Computing
Scientific computing systems are often complex and require high performance and accuracy. Defects in scientific computing can lead to incorrect results, degraded performance, or even compromised system security. Murphy et al. (2020) collect security defects from open-source scientific software projects written in Julia, a programming language designed for numerical analysis and computational science. Likewise, Azad et al. (2023) build a dataset by investigating 1,729 performance bugs from 23 real-world high-performance computing (HPC) applications such as molecular simulation and computational fluid dynamics. In addition to scientific computing applications, Franco et al. (2017) study numerical bugs across scientific computing libraries and present a dataset of 269 numerical bugs from six libraries, including NumPy (num, 2025), SciPy (sci, 2025), and LAPACK (lap, 2025). Collectively, these datasets provide valuable insights into the functional, performance, and security challenges faced in scientific computing.
@@ 3.1.8. Other Application Domains
In addition to the above, software defect datasets have been created for various other application domains. For example, for cloud computing, datasets of software defects are constructed from cloud systems (Gunawi et al., 2014) and cluster operators (Xu et al., 2024). For database-related software, datasets are available for both transaction bugs in database systems (Cui et al., 2024) and access bugs in database applications (Liu et al., 2024b). Other application domains include computational notebooks (de Santana et al., 2024) and internet of things (IoT) (Makhshari and Mesbah, 2021). While these domains have only a small number of datasets, these datasets provide insights into the challenges caused by domain-specific bugs.
@@ 3.2. Types of Defects
This section categorizes datasets based on the types of defects they contain. Across the 132 software defect datasets we survey, we identify four commonly considered defect types: functional bugs, security vulnerabilities, performance bottlenecks, and concurrency issues. In addition, we include a fifth category, others, to account for datasets that do not clearly fall into any of the four primary categories.
@@
Functional (Timperley et al., 2024; Andrade et al., 2024; Liu et al., 2024a, b; Zhang et al., 2024a; Cui et al., 2024; Yu et al., 2024; Xu et al., 2024; Waseem et al., 2024; Yu and Wang, 2024; Tian et al., 2024; Jimenez et al., 2024; Pramod et al., 2024; Silva et al., 2024; Song et al., 2023; Lee et al., 2024; An et al., 2023; Wu et al., 2024; Avula et al., 2023; Mahbub et al., 2023; Zhao et al., 2023a; Kim and Hong, 2023; Morovati et al., 2023; Saavedra et al., 2024; Jiang et al., 2023; Applis and Panichella, 2023; Guan et al., 2023; Xiong et al., 2023; Csuvik and Vidács, 2022; Pachouly et al., 2022; Liang et al., 2022; Keshavarz and Nagappan, 2022; Zhang et al., 2022b; Chen et al., 2022; Song et al., 2022; Paltenghi and Pradel, 2022; Liu et al., 2022; Wendland et al., 2021; Akimova et al., 2021; Gyimesi et al., 2021; Kim et al., 2021b; Ngo et al., 2021; Kamienski et al., 2021; Du et al., 2021; Makhshari and Mesbah, 2021; Ferenc et al., 2020; Zhang et al., 2020; Karampatsis and Sutton, 2020; Bures et al., 2020; Widyasari et al., 2020; Wang et al., 2020; Garcia et al., 2020; Vieira et al., 2019; Benton et al., 2019; Tomassi et al., 2019; Riganelli et al., 2019; Madeiral et al., 2019; Xu et al., 2019; Saha et al., 2018; Ferenc et al., 2018; Zhang et al., 2018; Madeyski and Kawalerowicz, 2017; Tan et al., 2017; Ohira et al., 2015; Le Goues et al., 2015, 2015; Just et al., 2014; Lamkanfi et al., 2013; D’Ambros et al., 2012; Gunawi et al., 2014; Ho et al., 2023; de Santana et al., 2024; Wang et al., 2023a, b; Zhao et al., 2023b; Wang et al., 2023c, 2022b; Shen et al., 2021; Wang et al., 2021a; Eghbali and Pradel, 2020; Gu et al., 2019; Franco et al., 2017; Lin et al., 2017; Ye et al., 2023a; Tan et al., 2018; Lu et al., 2005; Cifuentes et al., 2009) 87
Security (Ruan et al., 2024; Ni et al., 2024; Tebib et al., 2024; Jiang et al., 2024; Yu et al., 2024; Waseem et al., 2024; Das et al., 2024; Yu and Wang, 2024; Pramod et al., 2024; Cai et al., 2024; Wang et al., 2024; Zhong et al., 2023; An et al., 2023; Chen et al., 2023b; Senanayake et al., 2023; Brust et al., 2023; da Ponte et al., 2023; Bhuiyan et al., 2023; Zhang et al., 2022a; Bui et al., 2022; Pereira et al., 2022; Chen et al., 2022; Gyimesi et al., 2021; Paul et al., 2021; Zheng et al., 2021; Nikitopoulos et al., 2021; Afrose et al., 2023; Makhshari and Mesbah, 2021; Wu et al., 2020; Zhang et al., 2020; Murphy et al., 2020; Fan et al., 2020; Garcia et al., 2020; Radu and Nadi, 2019; Ponta et al., 2019; Lin et al., 2019; Gkortzis et al., 2018; Mitra and Ranganath, 2017; Ohira et al., 2015; Ku et al., 2007; Gunawi et al., 2014; Xu et al., 2023; Wang et al., 2023b) 43
Performance (Liu et al., 2024b; Cui et al., 2024; Yu et al., 2024; Waseem et al., 2024; Yu and Wang, 2024; Pramod et al., 2024; Kim and Hong, 2023; Guan et al., 2023; Kim et al., 2022; Gyimesi et al., 2021; Du et al., 2021; Makhshari and Mesbah, 2021; Sánchez et al., 2020; Zhang et al., 2020; Wang et al., 2020; Radu and Nadi, 2019; Zhang et al., 2018; Ohira et al., 2015; Zaman et al., 2012; Selakovic and Pradel, 2017; Gunawi et al., 2014; Ho et al., 2023; Azad et al., 2023; Zhao et al., 2023b; Wang et al., 2023c; Shen et al., 2021; Wang et al., 2021a; Franco et al., 2017; Lu et al., 2005) 29
Concurrency (Timperley et al., 2024; Liang et al., 2023; Yuan et al., 2021; Bagherzadeh et al., 2020; Gao et al., 2018; Lin et al., 2015; Jalbert et al., 2011; Gunawi et al., 2014; Shen et al., 2021; Wang et al., 2017; Lu et al., 2005) 11
Others (Liu et al., 2019; Chen et al., 2021b; Ren et al., 2020) 3
@@
Table 3 lists the datasets for each category, where a single dataset may appear under one or more categories. Functional defects are the most common type of defects in software defect datasets, with 87 (87/132) datasets in this category. Security vulnerabilities are the second most common type, which are present in 43 (43/132) datasets. Performance defects and concurrency issues are less common, with 29 (29/132) and 11 (11/132) datasets respectively. Figure 3 shows the breakdown of types of bugs for each of the application domains discussed in Section 3.1. We see that functional defects dominate most application domains and other types are also prevalent in specific domains based on the nature of the applications. The rest of this section further discusses these defect types and the datasets that target them. For each defect type, we discuss the characteristics of the defects and highlight representative datasets.

### v2

@@ 3.1.2. Defects in Compilers and Runtime Engines
Compilers and runtime engines are critical components of software systems, responsible for translating and executing code. Defects in compilers and runtime engines can significantly impact software performance, security, and execution correctness. We here mean by “runtime engine” both stand-alone implementations of a programming language, e.g., the Java Virtual Machine, and engines integrated into a larger software, e.g., a JavaScript engine embedded into a web browser. Datasets of defects in compilers and runtime engines often come as by-products of empirical research. For example, DBLP:conf/qrs/YuW24 conduct an empirical study on bugs in implementations of the Rust language , presenting a dataset of 194 bugs from Rust compilers and Rust infrastructure. DBLP:conf/uss/XuLDDLWPM23 investigate security bugs introduced by compilers for C/C++, highlighting the potential risks when silent bugs go undetected. In addition, as also mentioned above, DBLP:conf/issre/Du0MZ21 and DBLP:conf/sigsoft/ShenM0TCC21 study defects in deep-learning compilers, identifying common root causes and symptoms. For runtime engines, Bugs in Pods (DBLP:conf/issta/YuXZ0LS24) provides 429 bugs in the container runtime systems containerd , cri-o , gvisor , and runc . Additionally, DBLP:conf/icse/PaulTB21a provide a detailed dataset of security-related changes in the Chromium OS project, and DBLP:conf/msr/ZamanAH12 present a dataset of performance-related bugs in web browsers, focusing on how such bugs degrade system performance. Besides browsers, datasets constructed from studies of other runtime engines include defects in WebAssembly engines (DBLP:conf/wcre/WangZRLJ23; DBLP:journals/tosem/ZhangCWCLMMHL24), and bugs that impact the correctness and performance of JavaScript engines (DBLP:journals/infsof/WangBWYGS23). Altogether, these datasets are critical for understanding and addressing defects in compilers and runtime environments.
@@ 3.1.3. Defects in Mobile Applications
Mobile applications are prone to various types of defects, including functional bugs, security vulnerabilities, and resource issues. Several datasets, mostly based on Android applications, have been developed to represent these defects. For example, LVDAndro (DBLP:conf/secrypt/SenanayakeKAP023) and Ghera (DBLP:conf/promise/MitraR17) focus on security vulnerabilities, providing labelled vulnerability datasets and benchmarks for AI-based detection models and vulnerability analysis. Functional bugs are collected by Andror2 (DBLP:conf/msr/WendlandSMMHMRF21) and by DBLP:conf/issta/XiongX0SWWP0023, both of which analyze the behavior of defective mobile applications by manually reproducing the bugs. More specialized defects, such as data loss issues and resource leaks, are captured by DataLossRepository (DBLP:conf/msr/RiganelliMMM19) and DroidLeaks (DBLP:journals/ese/LiuWWXCWYZ19), respectively. Finally, DroixBench (DBLP:conf/icse/TanDGR18) provides defective APK files particularly for evaluating automated repair techniques on Android applications. These datasets play a crucial role in enhancing the reliability and security of Android applications by supporting bug detection, repair, and vulnerability research.
@@ 3.1.4. Defects Related to Blockchain
Blockchain, as a decentralized and distributed ledger technology, has been widely used in various applications, such as cryptocurrencies (Zheng et al., 2017). In general, blockchain applications are implemented as smart contracts, which are self-executing contracts with the terms of the agreement between multiple parties directly written into code. Defects in blockchain applications can have severe implications, such as financial loss. Several datasets focus on identifying and analyzing defects in smart contracts. From a general perspective, Jiuzhou (DBLP:conf/icsm/ZhangXL20) is a dataset of Ethereum smart contract bugs that are classified into 49 categories. Specifically for security vulnerabilities, SBP-Dataset (DBLP:journals/jss/CaiLZZS24) contains smart contract defects generated by removing guards and turning safe functions into unsafe ones. In addition, there are two datasets constructed from empirical studies. One defines 20 categories of smart contract bugs (DBLP:journals/tse/ChenXLGLC22), and the other identifies common patterns of bug fixes (DBLP:journals/jss/WangCHZBZ23). Unlike the above datasets, DBLP:conf/msr/KimKL22 focus on energy bugs in Ethereum client software instead of smart contracts. The dataset contains 507 bugs that may cause energy inefficiency mined from the buggy commits of client software operated in the Ethereum network.
3.1.5. Defects in Cloud Computing
Cloud computing systems provide on-demand access to computing resources over the internet and are prone to various software defects that can affect performance, reliability, and security. Datasets of software defects in cloud computing often arise from empirical studies. For instance, both DBLP:conf/cloud/GunawiHLPDAELLM14 and DBLP:conf/sigsoft/GaoDQGW0HZW18 investigate real-world issues in distributed cloud systems, with the latter focusing on crash recovery bugs. Additionally, DBLP:conf/issta/XuG024 present a dataset of bugs in cloud orchestration systems, while DBLP:journals/pacmpl/DrososSAM024 provide a dataset of misconfigurations in cloud infrastructure-as-code scripts.
@@ 3.1.6. Defects in Autonomous Systems
Autonomous systems, such as autonomous vehicles, aircraft, or robots, rely on robust software to operate safely and efficiently. Defects in autonomous systems can lead to severe consequences, such as accidents, injuries, and fatalities. DBLP:conf/sigsoft/WangLX0S21 and DBLP:conf/icse/GarciaF0AXC20 both present empirical studies of software defects in autonomous systems. The former study has collected a dataset of 569 bugs from PX4 and ArduPilot , which are well-known open-source autopilot software systems for unmanned aerial vehicles. The latter study presents a dataset of 499 autonomous vehicle bugs from Apollo and Autoware . In addition, ROBUST (DBLP:journals/ese/TimperleyHSDW24) offers a dataset of 221 bugs from seven popular robot operating systems. Together, these datasets highlight the unique challenges caused by defects in autonomous systems.
@@ 3.1.7. Defects Related to Quantum Computing
Defects in quantum computing systems, including both quantum applications and quantum platforms, present unique challenges due to the complexity and novelty of the technology. Bugs4Q (DBLP:journals/jss/ZhaoMLZ23) provides a benchmark of bugs in quantum programs, particularly focusing on the Qiskit framework , enabling controlled testing and debugging studies. Additionally, there are defect datasets constructed within empirical studies, such as one that gathers and categorizes bug patterns across various quantum platforms (DBLP:journals/pacmpl/PaltenghiP22), and another that focuses on defects within the intersection of quantum computing and machine learning (DBLP:conf/qsw/ZhaoWLLZ23). These datasets are useful for advancing the reliability of quantum software systems.
@@ 3.1.8. Defects Related to Scientific Computing
Scientific computing systems are often complex and require high performance and accuracy. Defects in scientific computing can lead to incorrect results, degraded performance, or even compromised system security. DBLP:conf/hotsos/MurphyBSR20 collect security defects from open-source scientific software projects written in Julia, a programming language designed for numerical analysis and computational science. Likewise, DBLP:conf/msr/AzadIHR23 build a dataset by investigating 1,729 performance bugs from 23 real-world high-performance computing (HPC) applications such as molecular simulation and computational fluid dynamics. In addition to scientific computing applications, DBLP:conf/kbse/FrancoGR17 study numerical bugs across scientific computing libraries and present a dataset of 269 numerical bugs from six libraries, including NumPy , SciPy , and LAPACK . Collectively, these datasets provide valuable insights into the functional, performance, and security challenges faced in scientific computing.
@@ 3.1.9. Other Application Domains
In addition to the above, software defect datasets have been created for various other application domains. For example, for database-related software, datasets are available for both transaction bugs in database systems (DBLP:conf/icse/CuiD0WSZWYXH0024) and access bugs in database applications (DBLP:journals/tosem/LiuMC24). Other application domains include computational notebooks (DBLP:journals/tosem/SantanaNAA24) and internet of things (IoT) (DBLP:conf/icse/Makhshari021). While these domains have only a small number of datasets, these datasets provide insights into the challenges caused by domain-specific bugs.
@@ 3.2. Types of Defects
Functional (DBLP:journals/ese/TimperleyHSDW24; DBLP:conf/edcc/AndradeLV24; DBLP:conf/msr/LiuHLZCSHM24; DBLP:journals/tosem/LiuMC24; DBLP:journals/tosem/ZhangCWCLMMHL24; DBLP:conf/icse/CuiD0WSZWYXH0024; DBLP:conf/issta/YuXZ0LS24; DBLP:conf/issta/XuG024; DBLP:conf/ease/WaseemDA0M24; DBLP:conf/qrs/YuW24; DBLP:conf/acl/TianYQCLPWHL0024; DBLP:conf/iclr/JimenezYWYPPN24; DBLP:conf/msr/PramodSTSW24; DBLP:conf/msr/SilvaSM24; DBLP:conf/kbse/SongWCCLLWP23; DBLP:conf/icst/LeeKYY24; DBLP:conf/kbse/AnKCYY23; DBLP:conf/sigsoft/WuLZ024; DBLP:conf/kbse/AvulaVM23; DBLP:conf/msr/MahbubSR23; DBLP:journals/jss/ZhaoMLZ23; DBLP:conf/icst/KimH23; DBLP:journals/ese/MorovatiNKJ23; DBLP:conf/icse/SaavedraSM24; DBLP:journals/tse/JiangLLZCNZHBZ23; DBLP:conf/msr/ApplisP23; DBLP:conf/icse/GuanXLLB23; DBLP:conf/issta/XiongX0SWWP0023; DBLP:conf/msr/CsuvikV22; DBLP:journals/softx/PachoulyAK22; DBLP:conf/icse/LiangLSSFD22; DBLP:conf/msr/KeshavarzN22; DBLP:conf/wcre/ZhangYYCYZ22; DBLP:journals/tse/ChenXLGLC22; DBLP:conf/issta/Song0NW0DM22; DBLP:journals/pacmpl/PaltenghiP22; DBLP:conf/issre/LiuZDHDM022; DBLP:conf/msr/WendlandSMMHMRF21; DBLP:conf/apsec/AkimovaBDKKMM21; DBLP:journals/stvr/GyimesiVSMBFM21; DBLP:conf/msr/KimKL21; DBLP:conf/splc/NgoNNV21; DBLP:conf/msr/KamienskiPBH21; DBLP:conf/issre/Du0MZ21; DBLP:conf/icse/Makhshari021; DBLP:journals/jss/FerencGGTG20; DBLP:conf/icsm/ZhangXL20; DBLP:conf/msr/KarampatsisS20; DBLP:conf/icst/BuresHA20; DBLP:conf/sigsoft/WidyasariSLQPTT20; DBLP:conf/msr/WangBJS20; DBLP:conf/icse/GarciaF0AXC20; DBLP:conf/promise/VieiraSRG19; DBLP:conf/icse/BentonGZ19; DBLP:conf/icse/DmeiriTWBLDVR19; DBLP:conf/msr/RiganelliMMM19; DBLP:conf/wcre/DelfimUMM19; DBLP:conf/dsa/XuYWA19; DBLP:conf/msr/SahaLLYP18; DBLP:conf/promise/FerencTLSG18; DBLP:conf/issta/ZhangCCXZ18; DBLP:conf/msr/MadeyskiK17; DBLP:conf/icse/TanYYMR17; DBLP:conf/msr/OhiraKYYMLFHIM15; DBLP:journals/tse/GouesHSBDFW15; DBLP:journals/tse/GouesHSBDFW15; DBLP:conf/issta/JustJE14; DBLP:conf/msr/LamkanfiPD13; DBLP:journals/ese/DAmbrosLR12; DBLP:conf/cloud/GunawiHLPDAELLM14; DBLP:conf/icsm/HoMI0SKNR23; DBLP:journals/tosem/SantanaNAA24; DBLP:journals/infsof/WangBWYGS23; DBLP:journals/jss/WangCHZBZ23; DBLP:conf/qsw/ZhaoWLLZ23; DBLP:conf/wcre/WangZRLJ23; DBLP:conf/kbse/WangWCCY22; DBLP:conf/sigsoft/ShenM0TCC21; DBLP:conf/sigsoft/WangLX0S21; DBLP:conf/kbse/EghbaliP20; DBLP:conf/compsac/GuWL0019; DBLP:conf/kbse/FrancoGR17; DBLP:conf/oopsla/LinKCS17; DBLP:conf/kbse/YeCG23; DBLP:conf/icse/TanDGR18; lu2005bugbench; BegBunch; DBLP:conf/icse/HumbatovaJBR0T20; DBLP:conf/kbse/RomanoLK021; DBLP:conf/msr/AmannNNNM16; DBLP:conf/msr/OliverDAH24; DBLP:conf/sigsoft/GaoDQGW0HZW18; DBLP:conf/sigsoft/Zhang0C0Z19; DBLP:conf/wcre/ReyesGSBM24; DBLP:conf/msr/BeyerGKRR24; DBLP:journals/access/AntalVKMHF24; DBLP:journals/ese/TambonNAKA24; DBLP:journals/pacmpl/DrososSAM024; DBLP:journals/tosem/TufanoWBPWP19; DBLP:conf/kbse/XuZL23) 100
Security (DBLP:conf/kbse/RuanLZL24; DBLP:conf/msr/NiSYZW24; DBLP:conf/pst/TebibAAG24; DBLP:journals/tosem/JiangJWMLZ24; DBLP:conf/issta/YuXZ0LS24; DBLP:conf/ease/WaseemDA0M24; DBLP:conf/kbse/DasAM24; DBLP:conf/qrs/YuW24; DBLP:conf/msr/PramodSTSW24; DBLP:journals/jss/CaiLZZS24; DBLP:conf/icse/WangHGWC024; DBLP:conf/kbse/ZhongLZGWGL23; DBLP:conf/kbse/AnKCYY23; DBLP:conf/raid/0001DACW23; DBLP:conf/secrypt/SenanayakeKAP023; DBLP:journals/compsec/BrustSG23; DBLP:conf/aina/PonteRM23a; DBLP:conf/icse/BhuiyanPVPS23; DBLP:conf/uss/ZhangP0W22; DBLP:conf/msr/BuiSF22; DBLP:conf/prdc/PereiraAV22; DBLP:journals/tse/ChenXLGLC22; DBLP:journals/stvr/GyimesiVSMBFM21; DBLP:conf/icse/PaulTB21a; DBLP:conf/icse/ZhengPLBEYLMS21; DBLP:conf/sigsoft/NikitopoulosDLM21; DBLP:journals/tse/AfroseXRMY23; DBLP:conf/icse/Makhshari021; DBLP:journals/jss/WuZCWM20; DBLP:conf/icsm/ZhangXL20; DBLP:conf/hotsos/MurphyBSR20; DBLP:conf/msr/FanL0N20; DBLP:conf/icse/GarciaF0AXC20; DBLP:conf/msr/RaduN19; DBLP:conf/msr/PontaPSBD19; DBLP:conf/icics/LinXZX19; DBLP:conf/msr/GkortzisMS18; DBLP:conf/promise/MitraR17; DBLP:conf/msr/OhiraKYYMLFHIM15; DBLP:conf/kbse/KuHCL07; DBLP:conf/cloud/GunawiHLPDAELLM14; DBLP:conf/uss/XuLDDLWPM23; DBLP:journals/jss/WangCHZBZ23; DBLP:conf/fie/SandersWA24; DBLP:conf/icse-apr/YadavW24; DBLP:conf/icse/HuZLYH024; DBLP:conf/issta/WuJPLD0BS23; DBLP:conf/sp/Dolan-GavittHKL16) 48
Performance (DBLP:journals/tosem/LiuMC24; DBLP:conf/icse/CuiD0WSZWYXH0024; DBLP:conf/issta/YuXZ0LS24; DBLP:conf/ease/WaseemDA0M24; DBLP:conf/qrs/YuW24; DBLP:conf/msr/PramodSTSW24; DBLP:conf/icst/KimH23; DBLP:conf/icse/GuanXLLB23; DBLP:conf/msr/KimKL22; DBLP:journals/stvr/GyimesiVSMBFM21; DBLP:conf/issre/Du0MZ21; DBLP:conf/icse/Makhshari021; DBLP:journals/access/SanchezDMS20; DBLP:conf/icsm/ZhangXL20; DBLP:conf/msr/WangBJS20; DBLP:conf/msr/RaduN19; DBLP:conf/issta/ZhangCCXZ18; DBLP:conf/msr/OhiraKYYMLFHIM15; DBLP:conf/msr/ZamanAH12; DBLP:conf/se/SelakovicP17; DBLP:conf/cloud/GunawiHLPDAELLM14; DBLP:conf/icsm/HoMI0SKNR23; DBLP:conf/msr/AzadIHR23; DBLP:conf/qsw/ZhaoWLLZ23; DBLP:conf/wcre/WangZRLJ23; DBLP:conf/sigsoft/ShenM0TCC21; DBLP:conf/sigsoft/WangLX0S21; DBLP:conf/kbse/FrancoGR17; lu2005bugbench; DBLP:conf/icse/HumbatovaJBR0T20; DBLP:journals/ese/TambonNAKA24; DBLP:journals/pacmpl/DrososSAM024; DBLP:journals/tse/ZhaoXBCL23) 33
Concurrency (DBLP:journals/ese/TimperleyHSDW24; DBLP:conf/asiaccs/LiangYDMHZ23; DBLP:conf/cgo/YuanLLLLX21; DBLP:journals/pacmpl/0001FSK20; DBLP:conf/msr/GaoYJLYZ18; DBLP:conf/kbse/LinMZCZ15; DBLP:conf/hotpar/JalbertPPS11; DBLP:conf/cloud/GunawiHLPDAELLM14; DBLP:conf/sigsoft/ShenM0TCC21; DBLP:conf/kbse/WangDGGQYW17; lu2005bugbench) 11
Others (DBLP:journals/ese/LiuWWXCWYZ19; DBLP:conf/icse/ChenYLCLWL21; DBLP:conf/iwpc/RenL020) 3
The ManyBugs and IntroClass Benchmarks (DBLP:journals/tse/GouesHSBDFW15) are counted as two distinct datasets. This convention applies to all other tables.
This section categorizes datasets based on the types of defects they contain. Across the 151 software defect datasets we survey, we identify four commonly considered defect types: functional bugs, security vulnerabilities, performance bottlenecks, and concurrency issues. In addition, we include a fifth category, others, to account for datasets that do not clearly fall into any of the four primary categories.
@@
Table 4 lists the datasets for each category, where a single dataset may appear under one or more categories. Functional defects are the most common type of defects in software defect datasets, with 100 (100/151) datasets in this category. Security vulnerabilities are the second most common type, which are present in 48 (48/151) datasets. Performance defects and concurrency issues are less common, with 33 (33/151) and 11 (11/151) datasets respectively. Figure 4 shows the breakdown of types of bugs for each of the application domains discussed in Section 3.1. We see that functional defects dominate most application domains and other types are also prevalent in specific domains based on the nature of the applications. The rest of this section further discusses these defect types and the datasets that target them. For each defect type, we discuss the characteristics of the defects and highlight representative datasets.